# AutoCrawler : A Progressive Understanding Web Agent for Web Crawler Generation

### v1

@@ 4.3 Generate with Golden Label
Table 3: Length of action sequence of AutoCrawler based on different LLMs in Swde dataset.
4.4 Further Study with AutoCrawler
The length of the action sequence is dependent on the LLMs capability.
To comprehensively explore the performance of different LLMs in understanding web page structure, we explore the impact of models on the number distribution of the steps. In particular, we collect all the action sequences and calculate the average steps of AutoCrawler with different LLMs. The experimental result is reported in Table 3.
We observe that AutoCrawler with stronger LLMs generate fewer lengths of action sequence. AutoCrawler with GPT4 generates 1.57 steps on average, while the AutoCrawler with Mistral 7B generates 3.82 steps on average. This phenomenon can be interpreted as more powerful models having a better understanding of the web page hierarchical structure, thus being able to accurately output the appropriate XPaths in longer/deeper web pages, thereby reducing the number of steps.
Figure 3: The curve on length and height compression ratio in Swde dataset.
The "U" curve of compression ratio
We define the length of HTML as the number of tokens in the HTML, and its height as the height of the DOM tree represented by the HTML. we define the compression ratio of length and height as the ratio of the length/height of the original web page to that of the web page after being pruned by AutoCrawler .
$Compression_{L}=\frac{\#\text{tokens of new HTML}}{\#\text{tokens of origin HTML}}$ (3)
$Compression_{H}=\frac{\#\text{height of new HTML}}{\#\text{height of origin HTML}}$
We calculate their compression ratio of the Correct case and rank LLMs based on their performance. Figure 3 shows the result. It is interesting to note that there is a "U" curve on both the length and height compression ratios. This phenomenon can be explained from two aspects: on one hand, when LLM is powerful, it can generate the correct XPath without the process of step-back to re-accessing the sub-DOM tree; on the other hand, when the model is weak, it is unable to effectively understand the hierarchical structure of web page, and thus cannot generate reliable, effective XPaths for the web page.
Good case Bad case
Case //div[@class=’gray200B-dyContent’]/ b[contains(text(),’Height:’)]/following-sibling::text() //div[@class=’infopage’]//h5[contains (text(), ’703-528-7809’)]
@@ 4.4 Further Study with AutoCrawler
XPath fragility within AutoCrawler
@@
We mainly focus on the fragility of text because these webpages are from the same websites (i.e. @class is a good characteristic for generating stable action sequences). Table 4 shows XPath expressions that rely on text. We aim to explore the reusability of generating XPath based on text features. We manually calculated the proportion of bad cases with two types of predicates, contains and equal  [^5] https://www.w3schools.com/xml/xpath_syntax.asp. The results in Table 5 show that the stronger LLMs capability, the lower the proportion of bad cases with AutoCrawler . However, it should be noted that the current SoTA LLM GPT-4 still suffers from an XPath fragility problem, which indicates that relying entirely on LLMs to generate reliable XPath still has some distance to go.
4.5 Error Analysis
We perform an analysis by looking at the recorded action sequence of AutoCrawler with GPT-4 and identify the following common failure modes. We mainly focus on the case that is categorized as unexecutable, over-estimate, and else.
Non-generalizability of webpages
The target information and corresponding webpage structures exhibit variations across different webpages, leading to a lack of generalizability in AutoCrawler (i.e., the inability to apply the same rules across all webpages in the same website) For instance, for the task "Please extract the name of the company that offers the job" in the website job-careerbuilder, most webpages contain the company name, but there is one webpage where the company name is listed as "Not Available" on another node of DOM tree.
Miss in multi-valued
Interesting Presented with the task of generating a crawler for extracting address in restaurant webpages or contact phone number from university websites, the target information is located in multiple locations in the webpage, such as the information bar, title, etc. Although AutoCrawler is capable of generating action sequences to extract portions of information, crafting a comprehensive action sequence that captures all of the information remains a challenge.
5 Related Work
5.1 Web Automation with LLMs
Many studies explore the concept of an open-world in web simulation environments Shi et al. (2017); Yao et al. (2023); Deng et al. (2023); Zhou et al. (2023), encompassing a broad spectrum of tasks found in real-life scenarios, such as online shopping, flight booking, and software development. Current web automation frameworks mainly aim to streamline the web environment Sridhar et al. (2023); Gur et al. (2023); Zheng et al. (2024) and to devise strategies for planning and interacting with the web Sodhi et al. (2023); Ma et al. (2023). However, these frameworks exhibit a lack of reusability, with agents heavily reliant on LLMs for even similar tasks, resulting in inefficiencies.
5.2 DOM-based Web Extraction
These methods utilize the hierarchical structure of the webpage. Method of this category includes rule-based Zheng et al. (2008), learning wrappers(i.e a DOM-specific parser that can extract content) Gulhane et al. (2011); Kushmerick (1997); Dalvi et al. (2011). Contemporary strategies employ distant supervision to autonomously create training samples by matching data from existing knowledge bases (KBs) with web sources Lockard et al. (2018, 2019). While this significantly lowers the effort required for annotation, it unavoidably leads to false negatives because of the incompleteness of knowledge bases (KBs) Xie et al. (2021).
6 Conclusion
In this paper, we introduce the crawler generation task and the paradigm that combines LLMs and crawlers to improve the reusability of the current web automation framework. We then propose AutoCrawler , a two-phase progressive understanding framework to generate a more stable and executable action sequence. Our comprehensive experiments demonstrate that AutoCrawler can outperform the state-of-the-art baseline in the crawler generation task.
Limitation
We introduce a paradigm that combines LLMs with crawlers for web crawler generation tasks and propose AutoCrawler to generate an executable action sequence with progressively understanding the HTML documents. Though experimental results show the effectiveness of our framework, there are still some limits to our work.
First, our framework is restricted to the paradigm in the information extraction task for vertical webpages. LLMs with crawlers provide high efficiency in open-world web IE tasks, but can hardly transfer to existing web environments such as Mind2Web Deng et al. (2023), WebArena Zhou et al. (2023).
Second, our framework rely on the performance of backbone LLMs. Enhancing LLMs’ ability to understand HTML is a very valuable research question, including corpus collection and training strategy. We will conduct research on HTML understanding enchancement in future work.
Ethic statement
We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
Use of Human Annotations
Risks
Appendix A Experiments
A.1 Main results on Extended Swde
Table D.2 shows the result. By comparing Table 3.1, we find that: 1) Under complex extraction task settings (multiple target values and ambiguous problem description), the closed-source LLMs perform better in generating executable action sequences compared to the open-source LLMs. 2) There are some tasks with unclear descriptions, such as the "Calendar System" and "Facilities and Programs Offered" on university websites, which affect the wrapper generation performance of all methods.
A.2 Main results on Ds1
Due to Ds1 only contains 166 hand-crafted webpages, and for each website, there are only two webpages, so we take one webpage for inference and the other for evaluation. Meanwhile, due to the number of the seed websites being equal to one, we test three methods without applying the synthesis module described in Section 3.3.
Table D.2 shows the result in the DS1 dataset. Among all LLMs with three methods, GPT4 + AutoCrawler achieves the best performance, and AutoCrawler beats the other two methods in all LLMs, which is consistent with the conclusion we make above.
Models Method Exec Eval IE Eval
Correct($\uparrow$) Unex.($\downarrow$) F1
GPT-3.5-Turbo COT 36.75 43.46 47.99
    - synthesis 27.56 57.24 34.44
[…]
    - synthesis 28.62 59.01 35.01
[…]
    - synthesis 44.52 29.33 58.44
Gemini Pro COT 29.69 47.19 41.81
    - synthesis 27.56 57.24 33.09
[…]
    - synthesis 28.62 59.01 37.60
[…]
    - synthesis 39.46 31.56 56.48
GPT4 COT 61.88 14.37 76.95
    - synthesis 46.88 30.00 61.20
[…]
    - synthesis 56.87 25.31 69.78
[…]
    - synthesis 65.31 11.87 80.41
Table 6: Ablation study on our two-stage framework. We report Correct, Unexecutable from the executive evaluation, and F1 score from the IE evaluation in Swde dataset.
Appendix B Analysis on AutoCrawler
B.1 Ablation Study
To further justify the effectiveness of each component of AutoCrawler , we perform an ablation study. The results are shown in Table A.2. It shows that: 1) AutoCrawlerwithout a second module still beat the other two baseline methods among different LLMs. 2) The second module of AutoCrawler , synthesis module, not only improves AutoCrawler , but also improves the performance of other methods. Using more webpages for inference can make the generated wrapper more stable and have better generalization.
Figure 4: The performance of GPT4 + AutoCrawler with different number of seed websites in Swde dataset.
B.2 Seed Websites
In all previous experiments, we fixed the number of seed websites $n_{s}=3$, which demonstrates the effectiveness of the synthesis module. In this experiment, we offer different numbers of seed webpages and test the performance of AutoCrawler with GPT4. The result is shown in Figure 4.
As the number of seed webpages increases, the correct ratio increases, while the unexecutable ratio decreases. It suggests that the performance of AutoCrawler can still be further improved by providing more seed webpages. In addition, the performance improvement reduces as the number increases, which shows that there is an upper limit to improve the performance of AutoCrawler by increasing the number of seed webpages.
B.3 Extraction Efficiency
Suppose the number of seed webpages is $n_{s}$, the number of webpages on the same website is $N_{\mathcal{W}}$, the time to generate a wrapper is $T_{g}$, the time to synthesis is $T_{s}$, and the time for extract information from a webpage with a wrapper is $T_{e}$. The total time for extracting all information from all websites with AutoCrawler is
$T_{1}=T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}$ (4)
$T_{2}=N_{\mathcal{W}}T_{d}$ (5)
In real-world scenario, there are many webpages from the same websites to be extracted. Although generating a wrapper takes more time than extracting directly from a single webpage, the extraction efficiency of subsequent webpages would be significantly improved. To explore how many webpages are needed to make AutoCrawler more efficient in web IE, we calculate the threshold of $N_{\mathcal{W}}$. Suppose $T_{1}\leq T_{2}$, we have
$T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}\leq N_{\mathcal{W}}T_{d}$ (6)
$N_{\mathcal{W}}\geq\frac{n_{s}T_{g}+T_{s}}{T_{d}-T_{e}}$ (7)
It should be noted that $T_{g}$ depends on $d_{max}$ in Algorithm 1 and can be roughly considered as $T_{g}\approx d_{max}T_{d}$. In our experimental settings, we set $d_{max}=5$ and $n_{s}=3$. Also, under the approximation that $T_{s}\approx T_{d}$ and $T_{d}\gg T_{e}$, AutoCrawler have better extraction efficiency when a website contains more than 16 webpages.
Model F1
Render-Full Hao et al. (2011) 84.30
FreeDOM Lin et al. (2020) 82.32
SimpDOM Zhou et al. (2021) 83.06
[…]
WebFormer Wang et al. (2022) 92.46
Reflexion + GPT4 82.40
AutoCrawler + GPT4 88.69
Table 7: Comparing the extraction performance (Precision, Recall and F1) of 5 baseline models to our method AutoCrawler using GPT4 on the Swde dataset. Each value of the supervised model in the table is trained on 1 seed site.
B.4 Comparison with supervised baselines
Table 7 shows the comparison with 5 baseline models in web information extraction on supervised learning scenarios. These models are trained on webpages in some seed websites and tested on the other websites. Although the comparison is unfair because our methods is in zero-shot settings, AutoCrawler beat most of them on F1 scores. It shows that by designing an appropriate framework, LLMs can surpass supervised learning methods in some web information extraction tasks.
Figure 5: Comparison with the other two baselines.
B.5 Comparison with COT & Reflexion
Figure 5 more intuitively shows the specific differences between different baselines in the experiment. The most significant difference between AutoCrawler and other methods lies in whether the hierarchical structure of web pages is utilized to help LLMs reduce the difficulty of complex web structures. COT only executes one turn while the other executes multiple turns and can learn from the failed execution of the wrapper. Compared to the Reflexion method, AutoCrawler employs top-down and step-back operations to prune the DOM tree during each XPath generation process, thereby reducing the length of the web page. In contrast, the Reflextion method can only reflect and regenerate after producing an unexecutable XPath, which does not effectively simplify the webpage.
@@ D.2 Module Prompt
Models Method Executable Evaluation IE Evaluation
Closed-source LLMs
GPT-3.5-Turbo COT 34.49 3.48 4.53 56.10 0.35 1.05 87.96 42.16 40.58
Reflexion 43.90 1.74 2.09 49.13 0.35 2.79 93.46 49.58 48.66
AutoCrawler 45.30 4.18 8.01 35.89 0.35 6.27 83.60 60.84 56.69
[…]
Reflexion 34.15 2.09 6.97 51.57 0.35 4.88 84.43 45.19 41.66
AutoCrawler 35.89 5.23 10.10 42.86 0.35 5.57 83.74 52.75 47.73
[…]
Reflexion 63.76 3.83 5.57 20.91 0.35 5.57 86.00 76.50 74.50
AutoCrawler 63.07 3.48 5.92 16.72 0.35 10.45 81.29 78.77 74.77
Open-source LLMs
CodeLlama COT 9.01 1.29 2.15 85.84 0.00 1.72 87.22 12.62 11.21
Reflexion 13.73 1.72 3.00 80.26 0.00 1.29 89.41 17.76 16.01
AutoCrawler 11.16 0.00 1.72 85.84 0.00 1.29 92.49 13.29 12.52
[…]
Reflexion 29.62 1.05 4.18 62.02 0.35 2.79 83.44 36.44 33.64
AutoCrawler 40.07 3.83 9.41 39.37 0.35 6.97 81.63 57.10 51.57
[…]
Reflexion 36.24 3.48 3.83 51.92 0.00 4.53 83.53 45.03 43.64
AutoCrawler 37.63 2.44 5.92 50.52 0.35 3.14 86.91 47.09 44.33
Models Method Executable Evaluation IE Evaluation
Closed-source LLMs
GPT-3.5-Turbo COT 32.65 4.08 8.16 53.06 0.00 2.04 90.56 43.54 41.16
Reflexion 36.73 8.16 4.08 51.02 0.00 0.00 95.56 44.22 43.75
AutoCrawler 48.98 4.08 0.00 44.90 0.00 2.04 94.90 51.70 52.38
[…]
Reflexion 20.25 10.13 1.27 65.82 0.00 2.53 88.83 26.93 27.66
AutoCrawler 43.04 15.19 3.80 34.18 0.00 3.80 93.76 55.97 56.92
[…]
Reflexion 50.60 10.84 4.82 33.73 0.00 0.00 96.85 62.65 63.50
AutoCrawler 57.83 15.66 4.82 16.87 0.00 4.82 92.88 74.95 75.52
Open-source LLMs
CodeLlama COT 2.70 2.70 5.41 89.19 0.00 0.00 78.72 10.62 9.19
Reflexion 8.82 0.00 5.88 85.29 0.00 0.00 94.12 14.41 12.69
AutoCrawler 13.51 0.00 5.41 81.08 0.00 0.00 84.12 18.92 17.39
[…]
Reflexion 22.78 6.33 1.27 69.62 0.00 0.00 94.15 28.03 28.20
AutoCrawler 36.71 11.39 6.33 43.04 0.00 2.53 91.59 48.52 48.23
[…]
Reflexion 22.89 6.02 3.61 65.06 0.00 2.41 90.21 31.43 32.04
AutoCrawler 39.76 10.84 6.02 42.17 0.00 1.20 90.43 51.39 50.28
Table 9: The executable evaluation and IE evaluation of LLMs with three frameworks in Ds1 dataset. We examine 6 LLMs, including 3 closed-source LLMs and 3 open-source LLMs.
Listing 1: Prompts for ZS_COT, Reflexion and AutoCrawler
Prompt-crawler (for COT, Reflexion and AutoCrawler):
Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
Instruction: {0}
Here are some hints:
1. Do not output the xpath with exact value or element appears in the HTML.
2. Do not output the xpath that indicate multi node with different value. It would be appreciate to use more @class to identify different node that may share the same xpath expression.
3. If the HTML code doesnt contain the suitable information match the instruction, keep the xpath attrs blank.
4. Avoid using some string function such as substring()’ and normalize-space()’ to normalize the text in the node.
Please output in the following Json format:
"thought": "", # a brief thought of how to confirm the value and generate the xpath
"value": "", # the value extracted from the HTML that match the instruction
"xpath": "", # the xpath to extract the value
Heres the HTML code:
{1}
Prompt-Reflexion (for Reflexion and AutoCrawler):
Heres the HTML extraction task:
Task description: Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
Instruction: {0}
We will offer some history about the thought and the extraction result. Please reflect on the history trajectory and adjust the xpath rule for better and more exact extraction. Heres some hints:
1. Judge whether the results in the history is consistent with the expected value. Please pay attention for the following case:
1) Whether the extraction result contains some elements that is irrelevent
2) Whether the crawler return a empty result
3) The raw values containing redundant separators is considered as consistent because we will postprocess it.
2. Re-thinking the expected value and how to find it depend on xpath code
3. Generate a new or keep the origin xpath depend on the judgement and thinking following the hints:
1. Do not output the xpath with exact value or element appears in the HTML.
2. Do not output the xpath that indicate multi node with different value. It would be appreciate to use more @class and [num] to identify different node that may share the same xpath expression.
3. If the HTML code doesnt contain the suitable information match the instruction, keep the xpath attrs blank.
Please output in the following json format:
"thought": "", # thought of why the xpaths in history are not work and how to adjust the xpath
"consistent": "", # whether the extracted result is consistent with the expected value, return yes/no directly
"value": "", # the value extracted from the HTML that match the task description
"xpath": "", # a new xpath that is different from the xpath in the following history if not consistent
And heres the history about the thought, xpath and result extracted by crawler.
{1}
Heres the HTML code:
{2}
Prompt-synthesis (for COT, Reflexion and AutoCrawler):
Youre a perfect discriminator which is good at HTML understanding as well. Following the instruction, there are some action sequence written from several HTML and the corresponding result extracted from several HTML. Please choose one that can be best potentially adapted to the same extraction task on other webpage in the same websites. Here are the instruction of the task:
Instructions: {0}
The action sequences and the corresponding extracted results with different sequence on different webpage are as follow:
{1}
Please output the best action sequence in the following Json format:
"thought": "" # brief thinking about which to choose
"number": "" # the best action sequence choosen from the candidates, starts from 0. If there is none, output 0.
Prompt-judgement (for AutoCrawler):
Your main task is to judge whether the extracted value is consistent with the expected value, which is recognized beforehand. Please pay attention for the following case:
1) If the extracted result contains some elements that is not in expected value, or contains empty value, it is not consistent.
2) The raw values containing redundant separators is considered as consistent because we can postprocess it.
The extracted value is: {0}
The expected value is: {1}
Please output your judgement in the following Json format:
"thought": "", # a brief thinking about whether the extracted value is consistent with the expected value
"judgement": "" # return yes/no directly
Prompt-stepback (for AutoCrawler):
Your main task is to judge whether the following HTML code contains all the expected value, which is recognized beforehand.
Instruction: {0}
And heres the value: {1}
The HTML code is as follow:
{2}
Please output your judgement in the following Json format:
"thought": "", # a brief thinking about whether the HTML code contains expected value
"judgement": "" # whether the HTML code contains all extracted value. Return yes/no directly.
@@ 3. Generate a new or keep the origin xpath depend on the judgement and thinking following the hints:
Table 13: Prompts for crawler generation task in Swde dataset.

### v2

@@ B.1 Comparison with COT & Reflexion
Table 10: Length of action sequence of AutoScraper based on different LLMs in Swde dataset.
Models 1 2 3 4 5 Avg.
GPT-4-Turbo 28 23 15 11 5 2.29
GPT-4-o-mini 50 10 5 1 16 2.06
GPT-3.5-Turbo 15 10 3 5 7 2.48
Gemini Pro 22 17 13 7 20 2.82
Mixtral 8$\times$7B 16 13 7 11 29 3.32
Phi-3-medium 14 15 6 2 46 3.61
Deepseek-coder 34 20 17 10 2 2.11
CodeLlama 18 6 6 9 33 3.46
Table 11: Length of action sequence of AutoScraper based on different LLMs in Ds1 dataset.
Models 1 2 3 4 5 Avg.
GPT-4-Turbo 61 40 45 53 76 3.15
GPT-4-o-mini 133 31 17 15 91 2.65
GPT-3.5-Turbo 88 35 48 23 97 3.02
Gemini Pro 60 41 29 28 132 3.45
Mixtral 8$\times$7B 51 38 26 29 138 3.59
Phi-3-medium 43 39 34 25 144 3.66
Deepseek-coder 120 79 35 33 20 2.14
CodeLlama 53 31 6 6 14 2.06
Table 12: Length of action sequence of AutoScraper based on different LLMs in Extended Swde dataset.
B.2 Further Study with AutoScraper
The length of the action sequence is dependent on the LLM capability.
To comprehensively explore the performance of different LLMs in understanding web page structure, we explore the impact of models on the number distribution of the steps. In particular, we collect all the action sequences and calculate the average steps of AutoScraper with different LLMs. The experimental result is reported in Table 1011 and 12.
We observe that AutoScraper with stronger LLMs generates fewer lengths of action sequence. AutoScraper with GPT-4-Turbo generates 1.57 steps on average, while the AutoScraper with Phi-3-medium generates 3.62 steps on average. This phenomenon can be interpreted as more powerful models having a better understanding of the web page hierarchical structure, thus being able to accurately output the appropriate XPaths in longer/deeper web pages, thereby reducing the number of steps.
@@ B.2 Further Study with AutoScraper
Good case Bad case
Case //div[@class=‘gray200B-dyContent‘]/ b[contains(text(),‘Height:‘)]/following-sibling::text() //div[@class=‘infopage‘]//h5[contains (text(), ‘703-528-7809‘)]
XPath fragility within AutoScraper
@@
We mainly focus on the fragility of text because these webpages are from the same websites (i.e. @class is a good characteristic for generating stable action sequences). Table 14 shows XPath expressions that rely on text. We aim to explore the reusability of generating XPath based on text features. We manually calculated the proportion of bad cases with two types of predicates, contains and equal  [^6] https://www.w3schools.com/xml/xpath_syntax.asp. The results in Table 13 show that the stronger LLMs capability, the lower the proportion of bad cases with AutoScraper . However, it should be noted that the current SoTA LLM GPT-4-Turbo still suffers from an XPath fragility problem, which indicates that relying entirely on LLMs to generate reliable XPath still has some distance to go.
@@ D.2 Module Prompt
Table 18: Prompts for scraper generation task in Swde dataset.
Prompt of Top-down Operation Heres the HTML extraction task: Task description: Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below. Instruction: {0} We will offer some history about the thought and the extraction result. Please reflect on the history trajectory and adjust the xpath rule for better and more exact extraction. Here are some hints: 1. Judge whether the results in the history are consistent with the expected value. Please pay attention to the following case: 1) Whether the extraction result contains some irrelevant elements 2) Whether the scraper returns an empty result 3) The raw values containing redundant separators are considered consistent because we will postprocess them. 2. Re-thinking the expected value and how to find it depends on the xpath code 3. Generate a new or keep the origin xpath depending on the judgement and thinking following the hints: 1. Do not output the xpath with the exact value or element that appears in the HTML. 2. Do not output the xpath that indicates multiple nodes with different values. It would be appreciated to use more @class and [num] to identify the different nodes that may share the same xpath expression. 3. If the HTML code doesnt contain suitable information to match the instruction, keep the xpath attribute blank. Please output in the following JSON format: { "thought": "", # thought of why the xpaths in history do not work and how to adjust the xpath "consistent": "", # whether the extracted result is consistent with the expected value, return yes/no directly "value": "", # the value extracted from the HTML that matches the task description "xpath": "", # a new XPath that is different from the XPath in the following history if not consistent } And heres the history of the thought, xpath and result extracted by scraper. {1} Heres the HTML code: ‘‘‘ {2} ‘‘‘
Prompt of Step-back Operation Your main task is to judge whether the following HTML code contains all the expected values, which are recognized beforehand. Instruction: {0} And heres the value: {1} The HTML code is as follows: ‘‘‘ {2} ‘‘‘ Please output your judgement in the following JSON format: { "thought": "", # a brief thinking about whether the HTML code contains expected value "judgement": "" # whether the HTML code contains all extracted value. Return yes/no directly. }
Prompt of Synthesis Youre a perfect discriminator who is good at HTML understanding as well. Following the instructions, there are some action sequences written from several HTML and the corresponding results extracted from several HTML. Please choose one that can be best potentially adapted to the same extraction task on other web pages on the same websites. Here are the instructions for the task: Instructions: {0} The action sequences and the corresponding extracted results with different sequences on different webpage are as follows: {1} Please output the best action sequence in the following JSON format: { "thought": "" # brief thinking about which to choose "number": " # the best action sequence chosen from the candidates, starts from 0. If there is none, output 0. }